determine the pH of a solution that is 0.15 M HClO2 (Ka=1.1 x 10^-2) and 0.15 M HClO (Ka=2.9 x 10^-8).
Answer #1

It is given that,

Concentration of HClO2 = 0.15 M,

Ka (HClO2) = 1.1 x 10-2

Concentration of HClO = 0.15 M,

Ka (HClO) = 2.9 x 10-8

Since, Ka (HClO2) = 1.1 x 10-2 which is greater than the Ka (HClO) = 2.9 x 10-8

So, HClO2 is a strong acid than the HClO. Then neglecting the HClO and considering HClO2.

Reaction is given by:-

HClO2 H+ + ClO2-

So, Ka = {[H+] x [ClO2-]} / HClO2 ........ (1)

Consider [H+] = [ClO2-] = X. Then HClO2 = 0.15 - X

Since, X is very small then (0.15 - X) will be almost equal to 0.15.

Then from equation (1), we have

Ka = X2/(0.15) = 1.1 x 10-2

or X2 = 0.15 x 1.1 x 10-2 = 0.165 x 10-2

or X = 0.041 M

So, [H+] = [ClO2-] = 0.041 M.

Then pH = -log(0.041) = 1.39

So, the pH of a solution will be, pH = 1.39